OpenStudy (anonymous):

how do you integrate: ∫ x/(x^2+y^2+z^2)^3/2 dx with respect to x

OpenStudy (anonymous):

\[1/2\int\limits_{?}^{?}d(x^2+y^2+z^2)/(x^2+y^2+z^2)\] 1/2ln(x^2+y^2+z^2)+C

OpenStudy (anonymous):

u substitution. u=x^2+y^2+z^2 du= 2x+2y+2z dx 1/2 du=x+y+z 1/2 ∫1/u^3/2 du = 1/2(-2/sqrt(u)) -1/sqrt(u) Substitute U back in and you have your answer

OpenStudy (anonymous):

that one makes sense. thank you!

OpenStudy (anonymous):

ugh, i sort of messed up on the du. du=2xdx, so 1/2 du= x dx.